OpenStudy (lucaz):

f(x)=-2x^2+400x is the function of a rectangular area, how can I find the largest area possible?

OpenStudy (loser66):

do you know derivative?

OpenStudy (lucaz):

without calculus, I think I need the vertex?

OpenStudy (loser66):

oh, we still have way to find out vertex, yes!! you are right

OpenStudy (loser66):

so? what is the vertex?

OpenStudy (lucaz):

hmm

OpenStudy (jdoe0001):

\(\bf f(x)=-2x^2+400x\implies f(x)=-2x^2+400x+0\\ \quad \\ \quad \\ f(x)=ax^2+bx+c\qquad \textit{vertex will be at }\left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (lucaz):

right, thanks

OpenStudy (loser66):

@lucaz I read your nick and realize that it is pronounced like "loca" by shakira. I forgot the name of the song until saw your nick, hihihi., enjoy http://www.youtube.com/watch?v=76VY054ghpk

OpenStudy (lucaz):

what a nice thing to know about my nick ¬¬